A copper rod of length 1.0 m is clamped at its middle point. Find the frequencies between 20 Hz and 20,000 Hz at which standing longitudinal waves can be set up in the rod. The speed of sound in copper is 3.8 km s-1.

Grade:11

1 Answers

Kevin Nash
9 years ago
Sol. a) Here given l = 5 cm = 5 × 10^–2 m, v = 340 m/s ⇒ n = V/2l = 340/2 * 5 * 10^-2 = 3.4 KHz ⇒then the highest harmonic in the audible range (20 Hz – 20 KHz) = 20000/3400 = 5.8 = 5 (integral multiple of 3.4 KHz).